1 kg of nitrogen (molar mass 28 kg/kmol) is compressed reversibly and isothermally from 1.01 bar , 20℃ to 4.2 bar. Calculate the work done and the heat flow during the process. Assume nitrogen to be a perfect gas.


1
Expert's answer
2022-03-15T16:49:21-0400

Solution.

"m=1kg;"

"M=28\sdot10^{-3} kg/mol;"

"P_1=1.01bar=101000Pa;"

"T=293K;"

"P_2=4.2bar=420000Pa;"

"A=\dfrac{m}{M}RTln\dfrac{V_2}{V_1};"

"P_1V_1=\dfrac{m}{M}RT\implies V_1=\dfrac{m}{M}\dfrac{RT}{P_1};"

"P_2V_2=\dfrac{m}{M}RT\implies V_2=\dfrac{m}{M}\dfrac{RT}{P_2};"

"\dfrac{V_2}{V_1}=\dfrac{P_1}{P_2};"

"A=\dfrac{1}{28\sdot10^{-3}}\sdot8.31\sdot293\sdot ln\dfrac{101000}{420000}=-124099J=-124kJ;"

Answer: "A=-124kJ;"

theat is rejected.